OpenStudy (anonymous):

a local dunkin doughnuts makes blueberry muffins that cost .69 each. past expericen shows that 15% of the muffins will spoil and have to be discarded. Assuming this dounut shop wants a 30% markup baed on cost and produces 200 muffins. each muffin should sale for?

OpenStudy (anonymous):

I got .90

OpenStudy (anonymous):

Based on 200 muffins, the total cost would be .69 * 200= 138 They want a markup of 30% based on costs, intended revenue = 138 *1.3 = 179.4 Since 15% of the muffins will spoil and be discarded. The no. of muffins left is, 0.85* 200 = 170 Price of the muffins, 179.4/170 = 1.06 approx.